# Tony Lindeberg

### v7

@@ 4.2 Relations to more detailed and fine-grained models of the primary visual cortex
Whereas one could possibly argue that the particle simulation approach, used for accumulating the receptive field histograms in Section 3.1.2 and 3.1.3, from explicitly generated samples from the distributions over the scale parameter ratio $\kappa$, could also be regarded as constituting numerical simulations with a certain numerical uncertainty, due to sampling issues of the underlying distributions, it would, in principle, also be possible to instead obtain more accurate estimates of the corresponding receptive field histograms without using any use of such sampled distributions. Instead, one can numerically solve the equations that determine the boundaries of the quantization bins in the distributions of the receptive fields, which, in turn, form the histograms as integrals of the underlying distributions over the resulting bin regions in the parameter space. Such an approach would then lead to results without any dependence on sampling issues regarding the model. In the experimental results reported in this paper, such possible sampling issues would nevertheless be expected to be very small, due to the rather large number of samples underlying the accumulation of the sampled histograms. Furthermore, the focus of this paper is on qualitative properties with regard to the distributions of receptive fields over different degrees of elongation.
@@
The proposed model does, from such a perspective, of judging whether the receptive field shapes could be regarded as spanning a substantial variability over the degree of elongation, constitute an as much simplified and idealized theoretical model as possible, while still respecting the covariance properties of the receptive fields under geometric image transformations. The model is, thus, also essentially parameter free, beyond the bounds on the scale parameter ratio $\kappa$, that were handled by symmetric intervals of the form $[1/\kappa_{\max},\kappa_{max}]$ for the here, for purposes of illustration only, arbitrary choice of $\kappa_{\max}=8$, and the above, for simplicity, in the absence of further information, assumed equal numbers of receptive fields corresponding to the different orders of spatial differentiation, in the composed histograms accumulated over receptive fields corresponding to multiple orders of spatial differentiation. Additionally, the underlying model for the visual receptive fields, which the presented work is based on, is theoretically very principled, by being derived in a purely axiomatic manner, and then after the axiomatic mathematical derivation a posteriori compared to actual neurophysiological recordings of biological receptive fields.
@@
A very interesting topic for future studies could nevertheless be to complement the presented theory with the influence of more complex neural mechanisms, such intra-V1-coupling mechanisms, the top-down influence in the visual cortex and inhibitory projections. Another interesting approach would be to integrate mechanisms for affine covariance into more fine-grained V1 models. Such more fine-grained modelling would then have the possibility to model detailed substructures in the reported orientation selectivity properties of the visual neurons, that are not comprised within the realm of our maximally idealized model. For example, Liu et al. (2015; 2017) have reported that the orientation tuning between the pinwheel and domain neurons in primary visual cortex may additionally depend also on the stimulus contrast and the stimulus size. [^4] Furthermore, an additional dependency on stimulus size could also result from an expansion of receptive field shapes over the size of the receptive fields, as also predicted from the property of affine covariance, that underlies the normative theory of visual receptive fields, used as the theoretical foundation for this study. From a theoretical viewpoint, such complementary , idependencies on stimulus contrast could be explained by the also non-linear dependencies of the receptive fields on the input stimuli, which may then lead to qualitatively different types of receptive fields from different stimuli, for example, based on mechanisms such as inhibitory tuning (Li et al. 2012) or, more generally, that the contrast of the visual stimuli may modulate the functional connectivity in the visual cortex (Nauhaus et al. 2009), see, however, also Alitto and Usrey (2004) and Nowak and Barone (2009) for complementary support regarding contrast invariant properties of the orientation tuning of the visual neurons.
@@
Author contributions
TL defined the scope of the work, developed the theory, performed the analysis, formulated the biological hypotheses, generated the illustrations in Figures 1, 2, 3, 4, 5, 8, 9, 10, 11, and 13, and wrote the paper.
Data availability statement
This paper does not contain any new experimental data, specifically neither on humans nor on cell cultures.
The data used for the analysis presented in this paper are based on existing sources in the scientific literature, explicitly cited in the text.
Additional information
The author declares no competing interests.
Appendix A Appendix
For purposes in the main article, we here extend the results concerning the orientation selectivity curves for idealized models of cells according to the generalized Gaussian derivative model for visual receptive fields in (Lindeberg 2024a) from first-order and second-order simple cells to also comprise third-order and fourth-order simple cells.
A.1 Derivations of orientation selectivity properties for third-order and fourth-order simple cells
@@
A.1.1 Third-order simple cell
Following the methodology in (Lindeberg 2024a) underlying the results summarized in Section 2.2.3, we will express an idealized model of a simple cell with four lobes along the preferred orientation of the simple cell as a third-order scale-normalized derivative of an affine Gaussian kernel (according to (5) for $m=3$), and for convenience of the calculations choose the preferred orientation as the horizontal $x_{1}$-direction (for $\varphi=0$) with spatial scale parameter $\sigma_{1}$ in the horizontal $x_{1}$-direction and spatial scale parameter $\sigma_{2}$ in the vertical $x_{2}$-direction, and thus with a spatial covariance matrix of the form $\Sigma_{0}=\operatorname{diag}(\sigma_{1}^{2},\sigma_{2}^{2})$:
@@ A.1.1 Third-order simple cell
A.1.2 Fourth-order simple cell
@@
A.1.3 Resulting orientation selectivity curves
@@
and with examples of graphs of these curves, for a few values of the scale ratio parameter $\kappa$, shown in Figure 5

### v10

@@ Appendix B Relations to more detailed and fine-grained models of the primary visual cortex
Whereas one could possibly argue that the particle simulation approach, used for accumulating the receptive field histograms in Section 4.1.2 and 4.1.3, from explicitly generated samples from the distributions over the scale parameter ratio $\kappa$, could also be regarded as constituting numerical simulations with a certain numerical uncertainty, due to sampling issues of the underlying distributions, it would, in principle, also be possible to instead obtain more accurate estimates of the corresponding receptive field histograms without using any use of such sampled distributions. Instead, one can numerically solve the equations that determine the boundaries of the quantization bins in the distributions of the receptive fields, which, in turn, form the histograms as integrals of the underlying distributions over the resulting bin regions in the parameter space. Such an approach would then lead to results without any dependence on sampling issues regarding the model. In the experimental results reported in this paper, such possible sampling issues would nevertheless be expected to be very small, due to the rather large number of samples underlying the accumulation of the sampled histograms. Furthermore, the focus of this paper is on qualitative properties with regard to the distributions of receptive fields over different degrees of elongation.
@@
If we consider the perspective of judging whether the receptive field shapes could be regarded as spanning a substantial variability over the degree of elongation. Then, the proposed model constitutes an as much simplified and idealized theoretical model as possible, while still respecting the covariance properties of the receptive fields under geometric image transformations. The model is, thus, also essentially parameter free, beyond the bounds on the scale parameter ratio $\kappa$, that were handled by symmetric intervals of the form $[1/\kappa_{\max},\kappa_{max}]$ for the here, for purposes of illustration only, arbitrary choice of $\kappa_{\max}=8$, and the above, for simplicity, in the absence of further information, assumed equal numbers of receptive fields corresponding to the different orders of spatial differentiation, in the composed histograms accumulated over receptive fields corresponding to multiple orders of spatial differentiation. Additionally, the underlying model for the visual receptive fields, which the presented work is based on, is theoretically very principled, by being derived in a purely axiomatic manner, and then after the axiomatic mathematical derivation a posteriori compared to actual neurophysiological recordings of biological receptive fields.
@@
A very interesting topic for future studies could nevertheless be to complement the presented theory with the influence of more complex neural mechanisms, such intra-V1-coupling mechanisms, the top-down influence in the visual cortex and inhibitory projections. Another interesting approach would be to integrate mechanisms for affine covariance into more fine-grained V1 models. Such more fine-grained modelling would then have the possibility to model detailed substructures in the reported orientation selectivity properties of the visual neurons, that are not comprised within the realm of our maximally idealized model. For example, Liu et al. (2015; 2017) have reported that the orientation tuning between the pinwheel and domain neurons in primary visual cortex may additionally depend also on the stimulus contrast and the stimulus size. [^10] Furthermore, an additional dependency on stimulus size could also result from an expansion of receptive field shapes over the size of the receptive fields, as also predicted from the property of affine covariance, that underlies the normative theory of visual receptive fields, used as the theoretical foundation for this study. From a theoretical viewpoint, such complementary dependencies on stimulus contrast could be explained by the also non-linear dependencies of the receptive fields on the input stimuli, which may then lead to qualitatively different types of receptive fields from different stimuli, for example, based on mechanisms such as inhibitory tuning (Li et al. 2012) or, more generally, that the contrast of the visual stimuli may modulate the functional connectivity in the visual cortex (Nauhaus et al. 2009), see, however, also Alitto and Usrey (2004) and Nowak and Barone (2009) for complementary support regarding contrast invariant properties of the orientation tuning of the visual neurons.
@@
Appendix C Derivations of orientation selectivity properties for third-order and fourth-order simple cells
For purposes in the main article, we here extend the results concerning the orientation selectivity curves for idealized models of cells according to the generalized Gaussian derivative model for visual receptive fields in (Lindeberg 2025) from first-order and second-order simple cells to also comprise third-order and fourth-order simple cells.
@@ Appendix C Derivations of orientation selectivity properties for third-order and fourth-order simple cells
C.1 Third-order simple cell
Following the methodology in (Lindeberg 2025) underlying the results summarized in Section 3.3.3 we will express an idealized model of a simple cell with four lobes along the preferred orientation of the simple cell as a third-order scale-normalized derivative of an affine Gaussian kernel (according to (5) for $m=3$), and for convenience of the calculations choose the preferred orientation as the horizontal $x_{1}$-direction (for $\varphi=0$) with spatial scale parameter $\sigma_{1}$ in the horizontal $x_{1}$-direction and spatial scale parameter $\sigma_{2}$ in the vertical $x_{2}$-direction, and thus with a spatial covariance matrix of the form $\Sigma_{0}=\operatorname{diag}(\sigma_{1}^{2},\sigma_{2}^{2})$:
@@ C.1 Third-order simple cell
C.2 Fourth-order simple cell
@@
C.3 Resulting orientation selectivity curves
@@
and with examples of graphs of these curves, for a few values of the scale ratio parameter $\kappa$, shown in in the bottom row of Figure 4.